# Debiased Sample Selection for Combating Noisy Labels

### v1

@@ Abstract
Learning with noisy labels aims to ensure model generalization given a label-corrupted training set. The sample selection strategy achieves promising performance by selecting a label-reliable subset for model training. In this paper, we empirically reveal that existing sample selection methods suffer from both data and training bias that are represented as imbalanced selected sets and accumulation errors in practice, respectively. However, only the training bias was handled in previous studies. To address this limitation, we propose a noIse-Tolerant Expert Model (ITEM) for debiased learning in sample selection. Specifically, to mitigate the training bias, we design a robust network architecture that integrates with multiple experts. Compared with the prevailing double-branch network, our network exhibits better performance of selection and prediction by ensembling these experts while training with fewer parameters. Meanwhile, to mitigate the data bias, we propose a mixed sampling strategy based on two weight-based data samplers. By training on the mixture of two class-discriminative mini-batches, the model mitigates the effect of the imbalanced training set while avoiding sparse representations that are easily caused by sampling strategies. Extensive experiments and analyses demonstrate the effectiveness of ITEM. Our code is available at this url ITEM.
@@ 1 Introduction
The remarkable generalization capability of deep neural networks (DNNs) is achieved through training on a large-scale dataset [1, 2]. However, existing training sets are usually collected by online queries [3], crowdsourcing [4], and manual annotations, which could inevitably incur wrong (or noisy) labels [5]. Since DNNs exhibit vulnerability to such low-quality datasets [6], training on the label-corrupted set presents a great challenge for modern DNNs. Therefore, to investigate learning with noisy labels (LNL) [7, 8, 9, 5] is practically important, which contributes to improvements of model’s generalization on real-world applications.
Figure 1: Existing selection criteria always lead to the imbalanced training set. We train ResNet-18 on noisy CIFAR-100 with two common selection strategies, including small-loss selection [10] and fluctuation-based noise filtering [11]. Class-level selected samples in the last epoch are counted.
Sample selection [10, 12, 11], a prevailing strategy for LNL, achieves considerable performance in mitigating the effects of noisy labels [13] by carefully selecting clean samples from the label-corrupted training set. The performance of sample selection approaches is largely decided by the selection criteria, which can be roughly categorized into two sides: 1) small-loss based strategies [10, 8, 14, 12, 15, 16], which are motivated by the memorization effect [17] that DNNs learn simple patterns shared by majority examples before fitting the noise. Hence, the samples with small losses in the early learning stage can normally be taken as clean samples. 2) fluctuation based strategies [18, 9, 11, 19], which are motivated by the observation that DNNs easily give inconsistent prediction results for noisy samples [11]. These methods normally consider an instance incorrectly labeled if its prediction results exhibit alterations within two consecutive training rounds or during a specified subsequent period.
Despite the validated effectiveness, there is a common view shared by existing sample selection methods that the sample selection processes are only influenced by a training bias, i.e., the accumulated error. In this paper, we for the first time show that by extensive experiments there exists another type of bias, i.e., the data bias, which is mainly caused by imbalanced data distribution. To be concrete, we split the bias in a selection-based learning framework into the training bias, which inherently exists in a self-training manner, and the data bias, which indicates the selected set tends to be class-imbalanced (see Figure 1). Therefore, the model is usually trained on class-imbalanced biased sets, resulting in sub-optimal results. Meanwhile, our experiments also provide some nontrivial findings, including 1) a corrupted selection tendency (see Figure 1(a)). The manner of self-training would exacerbate the imbalanced ratio in the training process. 2) class-discriminative selection performance (see Figure 1(b)). Although the overall selection performance is gradually improved as the training process proceeds, the performance on tail classes exhibits obvious degeneration. The above observations motivate us to consider a novel learning framework that can simultaneously deal with both training bias and data bias.
To this end, we propose a noIse-Tolerant Expert Model (ITEM), consisting of a noise-robust multi-experts structure and a mixed sampling strategy. Firstly, motivated by the idea of ensemble learning [20, 21], we design a robust network architecture that integrates the classifier with multiple experts to mitigate the training bias. Specifically, compared with the prevailing double-branch network [10, 12] where one network feeds selected samples to the other network, we independently conduct classifier learning and sample selection on different output layers. Therefore, incorrect selection results yielded by expert layers would not directly degenerate the classifier in subsequent learning, mitigating the accumulated error to a large degree. Meanwhile, compared with the double-branch network, our network is resource-friendly thanks to fewer trainable parameters and achieves better performance by combining multiple experts. Secondly, to confront the data bias, we propose a mixed sampling training strategy based on weighted data sampling. By calculating the quantity per class in the selected set, we obtain weights of different classes and further reverse weights via a mapping function for assigning larger weights to tail classes when sampling. Eventually, by using the mixup operation [22], our network is trained on a mixed mini-batch that combines two mini-batches respectively yielded by these two weighted samplers. This strategy fulfills class-balanced learning while avoiding sparse representations caused by the sampling strategy.
Our contribution can be summarized as follows:
  • We propose a noise-robust network architecture. Our multi-experts network is resource-friendly and exhibits improved robustness compared with the double-branch network. It shows significant potential in becoming the backbone to tackle noisy labels.
  • We propose a mixed sampling strategy, enabling the model to learn from tail classes via weighted samplers while avoiding sparse representations.
  • ITEM achieves significant performance on both synthetic and real-world noisy datasets, ensuring the effectiveness of our proposal on practical applications.
We also provide abundant analyses, verifying the effectiveness of each component in ITEM.
2 Related Work
Sample selection for LNL. Previous sample selection methods normally exploit the memorization effect of DNNs, i.e., DNNs first memorize training data with clean labels and then those with noisy labels [10, 23]. The small-loss criterion [10] is a typical method stemming from the memorization effect, which splits the noisy training set via a loss threshold. The samples with small losses are regarded as clean samples. Moreover, some researchers modified the small-loss criterion and proposed to select by leveraging the idea of disagreement [14], a regularization term (the Jensen-Shannon Divergence) [8], and a Mixture Gaussian Model for fitting the loss distribution [12], respectively. Besides, to mitigate the accumulated error caused by sample selection, current methods always resort to a double-branch network structure that ensembles selection results from two networks [10, 12, 8, 14, 24]. Besides, some studies [9, 18] introduce the uncertainty of loss to identify the noisy labels. i.e., predictions of noisy instances easily change and thus yield uncertain losses. A recent work [11] points out that hard samples are easily mixed with noisy samples and further discarded by the small-loss criterion. To solve this problem, Wei et al. [11] proposed a fluctuation-based sample selection method that considers the samples with inconsistent prediction results as incorrectly labeled data.
Long-tailed classification. Current long-tailed classification methods can be roughly categorized into two strategies, 1) customize the data sampler to modify sampling frequency for different classes (including over-sampling the tail-class samples and under-sampling the head-class samples) [25, 26, 27, 28], and 2) generate class-wise weights of training losses for improving the importance of tailed classes in terms of empirical risk minimization [29, 25, 28, 30, 31, 32]. However, both strategies result in underfitting in the representation of head classes, i.e., the model eventually learns sparse feature representations for head classes [33, 34]. To overcome this issue, some studies [35, 36] leverage the knowledge underlying head classes to promote the learning of tail classes. Cui et al. [37] designs a re-balance loss by computing the number of instances from each class and assigning different importance weights. Tang et al. [38] leverages causal learning and disentangles feature learning via backdoor adjustment.
(a) A corrupted selection tendency
(b) Class-level selection performance
@@ 2 Related Work
Figure 2: We train ResNet-18 on Sym. 60% CIFAR-100 with small-loss [10] and fluctuation-based noise filtering [11]. (a) Results on small-loss. Left: the class distribution of three training stages. Right: dynamic numbers of four representative categories. (b) Results on two selection criteria. Visualization of the class-level selection performance $\text{F}_{k}$ in 100- and 200-th epoch. The average F-score among all classes is reported. Although the performance increases, the change is different for varying classes.
Mixture-of-experts. MoE, which belongs to a combination method, has great potential to improve performance in machine learning [39]. Earlier works on MoE mainly focus on dividing the problem space between different experts. Concretely, combining classifiers exhibits promising performance in classification tasks [40, 41], especially for the problem that involves a limited number of patterns, high-dimensional feature sets, and highly overlapped classes [42, 43]. Nowadays, MoE has been applied to several fields, such as long-tailed classification [44, 45] and nature languages processing (NLP) [46, 47].
In this paper, we do not focus on the selection criterion but rather propose a new robust network architecture. Generally, our network can be widely applied to current sample selection methods and further improve their performance.
3 Background
@@
Assume $\mathcal{X}$ is the feature space and $\mathcal{Y}=\{1,2,\ldots,K\}$ is the label space. Suppose the training set is denoted by $D_{N}=\{({\bm{x}}_{i},y_{i})\}_{i\in[N]}$, where $[N]=\{1,2,\ldots,N\}$ is the set of indices. Since the annotator may give wrong labels in practice [48, 49], the learner thus can only observe a label-corrupted set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$.
As a prevailing strategy for LNL, sample selection [11, 50, 12, 9] aims to progressively select a reliable subset ${\tilde{D}_{M}}\subseteq{\tilde{D}_{N}}$ ($M<N$) and feed ${\tilde{D}_{M}}$ to the classifier for training. We introduce $v_{i}\in\{0,1\}$ to indicate whether the $i$-th instance is selected ($v_{i}=1$) or not ($v_{i}=0$). The performance of selection can be reflected by the deviation between the actual selected samples and the total clean samples, which can be measured by the F-score F, where $\text{F}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$, P and R denote selection precision and recall, respectively. In this paper, we individually calculate each class’s F-score for further analyses, i.e., $\text{F}_{k}$=$\frac{2\cdot\text{P}_{k}\cdot\text{R}_{k}}{\text{P}_{k}+\text{R}_{k}}$, $\text{P}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,\tilde{y}_{i}=k)}$, and $\text{R}_{k}$=$\frac{\sum\nolimits_{i\in[N]}\mathbbm{1}(v_{i}=1,y_{i}=\tilde{y}_{i}=k)}{\sum\nolimits_{i\in[N]}\mathbbm{1}(y_{i}=\tilde{y}_{i}=k)}$.
@@
  • Current selection criteria easily result in an imbalanced selection subset. As shown in Figure 1, we can observe that both the two adopted selection strategies incur the imbalanced data distribution under two noisy conditions. The reason is that both the two selection criteria rely on model performance. However, the model has different capacities for different classes. For those classes with indistinguishable characteristics (i.e., tail classes in the figure), the model tends to produce a large loss or inconsistent results and further discards those samples.
  • Sample selection, conducted as self-training, will exacerbate the imbalanced ratio during the training process. As shown in Figure 1(a), the imbalanced ratio of the selected set increases as the training process proceeds, i.e., the number of selected instances increases in head classes and decreases in tail classes (see the right part of Figure 1(a)). The intuitive reason is that the model’s performance on tail classes hardly improves due to the limited number of available samples for training, which further degrades the effectiveness of selection criteria on these classes.
  • The selection performance is inherently influenced by the imbalanced class distribution. As shown in Figure 1(b), the selection performance of both two selection criteria increases in head classes but decreases in tail classes. The reason is that the selected errors are relatively small for categories with higher F-scores, hence the training of these classes would further improve the model’s accuracy and selection performance. However, for tail classes with lower F-score, the self-training mechanism would further increase the selection error, and thus fail to correctly select instances in the follow-up phase.
Figure 3: Comparisons of various networks. (a) Typical classification network, which consists of a feature extractor and a classifier layer. (b) Mixture-of-experts (MoE) [51], a set of experts jointly gives the predicted label for the input. (c) Double-branch robust structure, the network is trained on a selected set which is considered as the clean by another network. (d) Ours, our framework consists of a classifier layer and a set of expert layers.
In LNL, we are given a label-corrupted training set ${\tilde{D}_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ of $N$ samples. A classifier with learnable parameters $f_{\bm{\theta}}$ is a function that maps from the input space $\mathcal{X}$ to the label space $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$. In multi-class classification, we always update the parameter $\bm{\theta}$ by minimizing the following empirical risk:
@@
where $\mathcal{L}(\cdot)$ is the given loss function, e.g., Softmax Cross-Entropy (CE) loss. In sample selection strategy, previous works mainly focus on designing high-effectively selection criteria. A universal training objective on the selected clean set can be written as
$\small\hat{R}_{\text{clean}}(f)$ $=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(f,\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]},\quad\text{where}\,\,\textsf{Criterion}(\cdot)=\left\{\begin{array}[]{lr}1,\,\,\text{If selected},&\\ 0.&\end{array}\right.$ (4)
We empirically show that directly optimizing ${\hat{R}}_{\text{clean}}(f)$ with an existing selection criterion is always biased (in Sec. 3), further resulting in inferior performance. For this, we propose to fulfill debiased learning by handling both the training and data bias. First, to solve the training bias, we design a novel network architecture, called noIse-Tolreant Expert Model (ITEM). Compared with the widely applied double-branch network, our proposal exhibits greater robustness to noisy labels. Second, to solve the data bias, we propose a mixed sampling strategy that can mitigate the side-effect caused by the class-imbalanced set $\tilde{D}_{\text{clean}}$ while avoiding sparse representations.
@@ 3.1 ITEM: noIse-Tolerant Expert Model
The training bias of sample selection stems from the way of a self-training manner. Previous works [10, 12] always resort to the double-branch network to confront this problem. Concretely, the network is trained on the clean set that is selected by the other network. Motivated by Mixture-of-experts [41, 51], we propose a robust architecture that integrates multiple experts into the classifier, which independently conducts classifier learning and sample selection in a single network.
@@
Specifically, compared with a normal classification network with a classifier layer $f$, our network ITEM contains $f$ and a set of expert layers $\{g^{1},...,g^{m}\}$ with size $m$. For robust selection, we propose to conduct existing criteria on expert layers. Therefore, the training objective in our framework can be summarized as
@@
$\small\hat{R}_{\text{clean}}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\textsf{Criterion}(\{g^{1},...,g^{m}\},\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$
Intuitively, the independent selection phase avoids the classifier layer selecting noisy samples and then updates its parameters on these data, which contributes to mitigating the training bias. We show the comparison of varying networks in Figure 3. Compared with the prevailing double-branch network, which is proved to be noise-robust in previous works [12, 10], our architecture contains two merits, including 1) resource-friendly training, in contrast to the feature extractor’s huge parameters, the last layer’s parameters for classification are significantly fewer. For example, the parameters of the classifier (or the expert) layer are less than 0.01 million, while the total parameters of a single ResNet-18 for CIFAR-10 are roughly 11.1 million. Thus, our network is more friendly to GPU and computer memory. 2) Greater performance, thanks to the property of fewer parameters, our model can integrate more experts, benefiting both selection and test accuracy by ensembling the results.

### v3

# Understanding and Mitigating the Bias in Sample Selection for Learning with Noisy Labels

@@ Abstract
Learning with noisy labels aims to ensure model generalization given a label-corrupted training set. The sample selection strategy achieves promising performance by selecting a label-reliable subset for model training. In this paper, we empirically reveal that existing sample selection methods suffer from both data and training bias, which are represented in practice as imbalanced selected sets and accumulation errors. However, only the training bias was handled in previous studies. To address this limitation, we propose a noIse-Tolerant Expert Model (ITEM) for debiased learning in sample selection. Specifically, to mitigate the training bias, we design a robust Mixture-of-Expert network that conducts selection and learning on different layers. Compared with the prevailing double-branch network, our network performs better on both selection and prediction by ensembling multiple experts while training with fewer parameters. Meanwhile, to mitigate the data bias, we propose a weighted sampling strategy that assigns larger sampling weights to classes with smaller frequencies. Using MixUp, the model is trained on a mixture of two batches: one sampled by a weighted sampler and the other by a regular sampler, which mitigates the effect of the imbalanced training set while avoiding sparse representations that are easily caused by sampling strategies. Extensive experiments on seven noisy benchmarks and analyses demonstrate the effectiveness of ITEM. The code is released at https://github.com/1998v7/ITEM.
@@ 1 Introduction
The remarkable generalization capability of deep neural networks (DNNs) is achieved through training on a large-scale dataset. However, existing training sets are usually collected by online queries [6], crowdsourcing [48], and manual annotations, which could inevitably incur wrong (or noisy) labels [40]. Since DNNs exhibit vulnerability to such low-quality training sets [52], training on the label-corrupted set presents a great challenge for the modern machine-learning community. To mitigate the adverse effect brought by noisy samples, learning with noisy labels (LNL) [37, 44] is important, contributing to improvements of the model’s generalization on practical applications.
Figure 1: Existing selection criteria always lead to an imbalanced training set, termed as the data bias. A ResNet-34 is trained on CIFAR-100N. We visualize the class distribution of the selected set, given three typical selection criteria. The quantity of class-level samples in the last epoch is counted, while the index of classes is sorted. More results can be found in Appendix A.
@@
Figure 2: We train ResNet-18 on Sym. 60% CIFAR-100 with small-loss [11] and fluctuation-based noise filtering [39]. (a) Results on small-loss. Left: the class distribution of three training stages. Right: dynamic numbers of four representative categories. (b) Results on two selection criteria. Visualization of the class-level selection performance (F-score) $\text{F}_{k}$ in 100- and 200-th epoch.
Sample selection [11, 21, 39], a prevailing strategy for LNL, achieves considerable performance in mitigating the effects of noisy labels [47] by carefully selecting clean samples from the label-corrupted training set. The performance of sample selection approaches is largely decided by the selection criteria, which can be roughly categorized into two sides: 1) small-loss based strategies [37, 11, 21, 50, 26, 3], which are motivated by the memorization effect [2] that DNNs learn simple patterns shared by majority examples before fitting the noise. Hence, the samples with small losses in the early learning stage can normally be taken as clean samples. 2) fluctuation based strategies [56, 39, 51, 44], which are motivated by the observation that DNNs easily give inconsistent prediction results for noisy samples. These methods [39, 51] normally consider an instance incorrectly labeled if its prediction results exhibit alterations within two consecutive training rounds or during a specified subsequent period.
Despite the validated effectiveness, there is a common view shared by existing sample selection methods that the sample selection processes are only influenced by a training bias, i.e., the accumulated error. In this paper, we, for the first time, show that by extensive experiments, there exists another type of bias, i.e., the data bias, which is mainly caused by imbalanced data distribution. To be concrete, we split the bias in a selection-based learning framework into the training bias, which inherently exists in a self-training manner, and the data bias, which indicates the selected set tends to be class-imbalanced (see Figure 1). Moreover, we customize a class-level F-score and do extensive experiments in the next section, providing three nontrivial findings that reveal the adverse effect of the data bias. These observations motivate us to consider a novel selection-based learning framework that can simultaneously address training and data bias.
To this end, we propose a noIse-Tolerant Expert Model (ITEM), consisting of a noise-robust multi-experts structure and a weighted sampling strategy.
  • Firstly, motivated by the idea of ensemble learning [9] and the Mixture-of-Expert model [16], we design a robust network structure integrating the classifier with multiple experts. Our network is more robust compared with existing robust networks like double-branch network [11] for the following reasons, 1) in our structure, the classifier learning and selection are disentangled, with experts selecting clean samples for the classifier’s training, naturally mitigating the issue of error accumulation. 2) Ensembling selection results from different experts reduces harmful interference from incompatible patterns (e.g., poor selection made by individual experts), leading to a cleaner selected set.
  • Secondly, to mitigate the implicit data bias during selection, we propose a weighted resampling strategy named mixed sampling. By calculating the quantity of each class in the selected set, we assign larger weights to tail classes through a mapping function during sampling. Finanlly, introducing the MixUp [53], our network is trained on a mixed batch that combines two batches based on a regular and a weighted sampler, fulfilling class-balanced learning while avoiding the issue of sparse representations.
The promising performance of ITEM is verified on seven noisy benchmarks. Extensive ablation studies and analyses also demonstrate the effectiveness of each component.
2 Understanding the Bias in Sample Selection
@@
Assume $\mathcal{X}$ is the feature space and $\mathcal{Y}=\{1,2,\ldots,K\}$ is the label space. Suppose the training set is denoted by ${\tilde{D}}=\{({\bm{x}}_{i},y_{i})\}_{i\in[N]}$, where $[N]=\{1,2,\ldots,N\}$ is the set of indices. Since the annotator may give wrong labels in practice [18], the learner thus can only observe a label-corrupted set ${D_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ with noisy labels.
As a prevailing strategy for LNL, sample selection [39, 24, 21, 44] aims to progressively select a reliable subset ${D_{M}}\subseteq{D_{N}}$ ($M<N$) and feed ${D_{M}}$ to the classifier for training. We introduce $v_{i}\in\{0,1\}$ to indicate whether the $i$-th instance is selected ($v_{i}=1$) or not ($v_{i}=0$). Generally, the performance of a selection-based method can be reflected by the deviation between the actual selected samples and the total clean samples, which can be measured by the F-score F, where $\text{F}=\frac{2\cdot\text{P}\cdot\text{R}}{\text{P}+\text{R}}$, P and R denote selection precision and recall, respectively. In this paper, we individually calculate each class’s F-score for further analysis. Specifically, the class-level F-score for class $k$ is written as
Fk=2PkRkPk+Rk,where
[…]
Figure 3: Comparisons of different architectures. (a) Typical classification network, which consists of a feature extractor and a classifier layer. (b) Mixture-of-experts (MoE) [29], a set of experts jointly gives the predicted label for the input. (c) Double-branch robust structure, the network is trained on a selected set that is considered clean by another network. (d) Ours, a mixture-of-experts module, is integrated into the classification network, which works for robust selection as well as prediction ensemble.
@@
  • Current selection criteria easily result in an imbalanced selection subset. As shown in Figure 1, we can see that both of the two adopted selection strategies incur the imbalanced data distribution under two noisy conditions. The reason is that both of the two selection criteria rely on model performance. However, the model has different capacities for different classes. For those classes with indistinguishable characteristics (i.e., rare classes), the model tends to produce a large loss or inconsistent results and further discards those samples.
  • Sample selection, conducted as self-training, will exacerbate the imbalanced ratio during the training process. As shown in Figure 2 (a), the imbalanced ratio of the selected set increases as the training process proceeds, i.e., the number of selected instances increases in head classes and decreases in tail classes (see the right part of Figure 2). The intuitive reason is that the model’s performance on tail classes hardly improves due to the limited number of available samples for training, which further degrades the effectiveness of selection criteria on these classes.
  • The selection performance is inherently influenced by the imbalanced class distribution. As shown in Figure 2 (b), the selection performance of both selection criteria increases in head classes but decreases in tail classes. The reason is that the selected errors are relatively small for categories with higher F-scores. Hence, the training of these classes would further improve the model’s accuracy and selection performance. However, for tail classes with lower F-score, the self-training mechanism would further increase the selection error and thus fail to select instances in the follow-up phase correctly.
Based on these observations, we can deduce that the classifier under a sample selection framework implicitly suffers from a data bias, which is, however, rarely studied in previous literature. Motivated by this, we aim to design a model that confronts not only accumulated error but also the imbalanced selected set.
3 Methodology
Preliminaries. In LNL, we are given a label-corrupted training set ${D_{N}}=\{({\bm{x}}_{i},{\tilde{y}}_{i})\}_{i\in[N]}$ of $N$ samples. A classifier with learnable parameters $f_{\bm{\theta}}$ is a function that maps from the input space $\mathcal{X}$ to the label space $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$. In multi-class classification, we always update the parameter $\bm{\theta}$ by minimizing the following empirical risk:
@@ 3 Methodology
where $\mathcal{L}(\cdot)$ is the given loss function, e.g., Softmax Cross-Entropy (CE) loss. In sample selection strategy, previous works mainly focus on designing high-effectively selection criteria. Here, we give a universal training objective on the selected clean set, which is written as
$\hat{R}_{\text{clean}}(f)$ $=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}(f,\bm{x}_{i},{\tilde{y}}_{i})\cdot\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]},\quad$
$\text{where}\,\,\mathrm{Criterion}(\cdot)=\left\{\begin{array}[]{lr}1,\,\,\text{If selected},&\\ 0.&\end{array}\right.$ (4)
In this unified formula, the classifier intended to be optimized is consistent with the classifier serving as the noise filter during the selection phase, which easily leads to the training bias. Empirical results in the section above show that directly optimizing ${\hat{R}}_{\text{clean}}(f)$ with an existing selection criterion is always biased, further resulting in inferior generalization performance.
Overview. We propose to fulfill debiased learning by handling both the training and data bias. First, to solve the training bias, we design a novel network architecture called noIse-Tolreant Expert Model (ITEM). Compared with the widely applied double-branch network, our proposal exhibits greater robustness to noisy labels. Second, to solve the data bias, we propose a weighted resampling strategy that can mitigate the side-effect caused by the class-imbalanced set $D_{\text{clean}}$ while avoiding sparse representations.
@@ 3.1 ITEM: noIse-Tolerant Expert Model
The training bias in sample selection arises from the self-training approach. Previous works [11, 21] always resort to the double-branch network to confront this problem. Concretely, the network is trained on the clean set that is selected by the other network. Motivated by Mixture-of-experts [31, 29], we propose a robust architecture that integrates multiple experts into the classifier, which independently conducts classification and sample selection.
@@
Specifically, compared with a normal classification network with a classifier layer $f$, our network ITEM contains $f$ and a set of additional expert layers $\{g^{1},...,g^{m}\}$ with size $m$. Each expert layer has the same dimension as the classifier layer, i.e., $\|f\|=\|g\|$. For robust selection, we propose to conduct existing criteria on expert layers instead of the classifier layer. Consequently, the training objective in our framework can be summarized as
@@
$\hat{R}_{\text{clean}}(f)=\frac{1}{N}\sum\nolimits_{i=1}^{N}\big{[}\mathrm{Criterion}$ ($(\{g^{1},...,g^{m}\},\bm{x}_{i},{\tilde{y}}_{i})\cdot$
$\mathcal{L}(f(\bm{x}_{i},\bm{\theta}),{\tilde{y}}_{i})\big{]}.$ (5)
Intuitively, the independent selection phase prevents the classifier layer from selecting noisy samples and updating its parameters based on them, significantly mitigating the training bias. In practice, a voting strategy is applied to ensemble the selection results of $m$ experts, where the sample selected by all experts is considered clean. It is noteworthy that the selection criterion $\mathrm{Criterion}(\cdot)$ is not restricted to a specific form (e.g., small-loss based or fluctuation-based), which highlights the applicability of this expert model as a backbone in sample selection.
Advantages of the MoE structure. We compare different network structures in LNL in Figure 3. Compared with the prevailing double-branch networks, which have proved to be noise-robust in previous works [11, 21], our architecture offers two advantages, including 1) Greater robustness to noisy labels. Each expert exhibits different capabilities during training, resulting in a better inductive bias for both selection and prediction. Specifically, by ensembling the selections from different experts, the MoE minimizes negative interference from incompatible patterns (e.g., poor selection made by individual experts), thereby enhancing the framework’s robustness against noisy labels. 2) Resource-friendly training. In contrast to the feature extractor’s huge parameters, the last layer’s parameters for classification are significantly fewer. The additional learnable parameters in our model are less than 2% of those in the double-branch network. Thus, our network is more friendly to GPU and computer memory.
3.2 Beta-Resampling